# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 4.2 Arrangements of tubes that cluster into rectangular prisms
It remains to show that Item (ii) holds for $\mathbb{T}^{\prime}$; as explained above it suffices to consider $\rho$-tubes with $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$, and convex sets $R$ that are $s\times t\times 1$ prisms, with $s\in[\delta,\rho]$ of the form $s=\delta^{j/N}$ and $t\in[s,\rho]$. We would like to estimate the number of tubes from $\mathbb{T}^{\prime}[T_{\rho}]$ contained in $R$.
By Conclusion (B) of Lemma 4.10, we have
@@
#$\#\{T\in\mathbb{T}^{\prime}[T_{\rho}]$ $\colon T\subset R\}\leq\delta^{-\eta_{2}}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}^{\prime}_{s})}\lesssim\delta^{-\eta_{2}-\varepsilon/4}ts(s^{\beta-2}\delta^{-\beta})$
$\lesssim\delta^{-\eta_{2}-\varepsilon/4}ts\rho^{\beta-2}\delta^{-\beta}=\delta^{-\eta_{2}-\varepsilon/4}\frac{ts}{\rho^{2}}\frac{\delta^{-\beta}}{\rho^{-\beta}}\leq\delta^{-2\eta_{2}-\varepsilon/4}\frac{ts}{\rho^{2}}(\#\mathbb{T}^{\prime\prime}[T_{\rho}]).$
Selecting $\eta_{2}$ sufficiently small, we obtain
#$\#\{T\in\mathbb{T}^{\prime}[T_{\rho}]\colon T\subset R\}\leq\delta^{-\varepsilon}\frac{|R|}{|T_{\rho}|}(\#\mathbb{T}^{\prime}[T_{\rho}]).$
@@
We conclude that $\mathbb{T}$ satisfies Condition (B) from Definition 2.15, and hence Conclusion (B) of Proposition 4.1 holds. ∎
5 Sticky Kakeya and the Self-Similar Wolff Axioms
@@ 5.1 Extremal Kakeya sets satisfy the Self-Similar Convex Wolff Axioms
Suppose $\omega$ from Definition 3.1 is positive. Then for all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be $\eta$ Assouad Extremal. Then there exists $\mathbb{T}^{\prime}\subset\mathbb{T}$ so that $(\mathbb{T}^{\prime},Y)_{\delta}$ is $\varepsilon$ Assouad extremal, and $\mathbb{T}^{\prime}$ satisfies the Self-Similar Convex Wolff axioms with error $\delta^{-\varepsilon}$.
@@
Let $\delta\in(0,\delta_{0}]$ and let $(\mathbb{T},Y)_{\delta}$ be $\eta$ Assouad Extremal. By pigeonholing, we can select a subset $\mathbb{T}_{1}\subset\mathbb{T}$ with #$\#\mathbb{T}_{1}\gtrsim\delta^{\eta}(\#\mathbb{T})$ so that $|Y(T)|\geq\frac{1}{2}\delta^{\eta}|T|$ for each $T\in\mathbb{T}_{1}$. Then #$\#\mathbb{T}_{1}=\delta^{-\beta}$ for some $\beta\geq\alpha-2\eta$ (recall that $\alpha$ is defined in (3.4)), and $\mathbb{T}_{1}$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}$. Furthermore, for all $\rho_{1},r_{1}$ with $\rho_{1}\leq\delta^{\eta}r_{1}$, and all balls $B$ of radius $r_{1}$, we have
@@
(In the above equation, $\omega(\beta+\tau)$ refers to the function $\omega$, evaluated at the (positive) number $\beta+\tau$). Provided $\eta_{1}$ is selected sufficiently small (depending on $\tau$), we have $\beta+\tau\geq\alpha-2\eta_{1}+\tau>\alpha$, and hence by the definition of $\alpha$ from (3.4), we have $\gamma>0$. Hence if $\eta_{1}$ is selected sufficiently small, then
@@
Step 4. We will show that Conclusion (D) from Proposition 4.1 cannot hold. Suppose to the contrary that Conclusion (D) holds, i.e. there exists $\rho\in[\delta^{1-\varepsilon},1]$, a $\rho$-tube $T_{\rho}$, and a set $\tilde{\mathbb{T}}\subset\mathbb{T}_{1}[T_{\rho}]$ satisfying Items (D.i) and (D.ii) from Conclusion (D). Let $\tilde{\delta}=\delta/\rho\in[\delta,\delta^{\varepsilon}]$, and let $\tilde{\mathbb{T}}_{\tilde{\delta}}$ be unit rescaling of $\tilde{\mathbb{T}}$ relative to $T_{\rho}$. Let $\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})$ be the image of $\tilde{Y}(\tilde{T})$ under this same rescaling. We have $|\tilde{Y}_{\tilde{\delta}}(\tilde{T})|\gtrsim\delta^{\eta}|\tilde{T}_{\tilde{\delta}}|\geq\tilde{\delta}^{\eta/\varepsilon}|\tilde{T}_{\tilde{\delta}}|$, and we will select $\eta$ and $\delta_{0}$ sufficiently small so that $|\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})|\geq\tilde{\delta}^{\eta_{1}}|\tilde{T}_{\tilde{\delta}}|$ for all $\tilde{T}_{\tilde{\delta}}\in\tilde{\mathbb{T}}_{\tilde{\delta}}$. The tubes in $\tilde{\mathbb{T}}_{\tilde{\delta}}$ are essentially distinct; by Item (D.i) we have #$\#\tilde{\mathbb{T}}_{\tilde{\delta}}\geq\tilde{\delta}^{-\beta-\tau}$, and by Item (D.ii) we have that $\tilde{\mathbb{T}}_{\tilde{\delta}}$ satisfies the Convex Wolff Axioms with error $\lesssim\tilde{\delta}^{-\eta_{1}}$.
@@
$\Big{|}B_{2}\cap N_{\rho_{2}}\Big{(}\bigcup_{\tilde{T}_{\tilde{\delta}}\in\tilde{\mathbb{T}}_{\tilde{\delta}}}\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})\Big{)}\Big{|}\geq(\rho_{2}/r_{2})^{\omega-\gamma/8}|B_{2}|.$
Undoing the unit rescaling described above, we conclude there exists an ellipsoid $E$ of dimensions $\rho r_{2}\times\rho r_{2}\times\rho$ so that
$\Big{|}E\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}\geq(\rho_{2}/r_{2})^{\omega-\gamma/8}|E|.$
Indeed, the above inequality follows from the observation that the image of a $\rho_{2}$-ball centered at $x$ under the above rescaling is an ellipse of dimensions $\rho\rho_{2}\times\rho\rho_{2}\times\rho_{2}$, and this ellipse contains a $\rho\rho_{2}$-ball centered at $x$. In particular, we can find a $\rho r_{2}$-ball $B$ with $|B\cap E|\sim|B|$, so that
$\Big{|}B\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}\gtrsim(\rho_{2}/r_{2})^{\omega-\gamma/8}|B|.$ (5.5)
Finally, observe that $\frac{\rho\rho_{2}}{\rho r_{2}}=\frac{\rho_{2}}{r_{2}}\leq\rho^{\eta_{1}}\leq\delta^{\eta_{1}\varepsilon}\leq\delta^{\eta}$ (provided we select $\eta\leq\eta_{1}\varepsilon$). Thus if we choose $\eta\leq\gamma/9$ and $\delta_{0}$ sufficiently small, then (5.5) contradicts (5.1). Hence Conclusion (D) cannot hold.
@@
It suffices to show that $\omega=0$. Suppose instead that $\omega>0$. Let $\eta_{0},\delta_{0}$ be the output from Theorem 5.2, with $\varepsilon=\omega/3$. By Proposition 5.1, there exists $\delta\in(0,\delta_{0}]$ and a set $(\mathbb{T}^{\prime},Y)_{\delta}$ with discretized Assouad dimension $\leq 3-\omega/3$ that satisfies the Self-Similar Convex Wolff Axioms (and hence the Convex Wolff Axioms at every scale) with error $\delta^{-\eta_{0}}$, and for which $\{Y(T)\}$ is a $\delta^{\eta_{0}}$ dense shading. But then (6.2) says that $\bigcup Y(T)$ has discretized Assouad dimension at least $33-\omega/$ 3, which is a contradiction. ∎
@@
6 Kakeya estimates for tubes satisfying the Convex Wolff Axioms at all scales

### v2

@@ 4.2 Arrangements of tubes that cluster into rectangular prisms
It remains to show that Item (ii) holds for $\mathbb{T}^{\prime}$. It suffices to consider $\rho$-tubes $T_{\rho}$ with $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$. Initially, we will consider convex sets $R$ that are $s\times t\times 1$ prisms, with $s\in[\delta,\rho]$ of the form $s=\delta^{j/N}$ and $t\in[s,\rho]$. We would like to estimate the number of tubes from $\mathbb{T}^{\prime}[T_{\rho}]$ contained in $R$.
By Conclusion (B) of Lemma 4.9 and (4.38), we have
@@
#$\#\{T\in\mathbb{T}^{\prime}[T_{\rho}]\colon T\subset R\}$ $\leq\delta^{-\eta_{2}}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}_{s})}\lesssim\delta^{-\eta_{2}-\varepsilon/4}(ts)(s^{\beta-2}\delta^{-\beta})\lesssim\delta^{-\eta_{2}-\varepsilon/4}(ts)(\rho^{\beta-2}\delta^{-\beta})$
$=\delta^{-\eta_{2}-\varepsilon/4}\Big{(}\frac{ts}{\rho^{2}}\Big{)}\Big{(}\frac{\rho}{\delta}\Big{)}^{\beta}\leq\delta^{-2\eta_{2}-\varepsilon/4}\Big{(}\frac{ts}{\rho^{2}}\Big{)}\big{(}\#\mathbb{T}^{\prime}[T_{\rho}]\big{)}\leq\delta^{-\varepsilon/2}\frac{|R|}{|T_{\rho}|}\big{(}\#\mathbb{T}^{\prime}[T_{\rho}]\big{)}.$
Finally, every convex set $W\subset T_{\rho}$ can be contained in a bounded number of $s\times t\times 1$ rectangular prisms $R$, where $s$ is of the form $\delta^{j/N}$, $t\in[s,\rho]$, and $|R|\leq\delta^{-\eta_{2}}|W|$. Thus for every convex set $W\subset\mathbb{T}_{\rho}$, we have
#$\#\{T\in\mathbb{T}^{\prime}[T_{\rho}]\colon T\subset W\}\leq\delta^{-\varepsilon}\frac{|W|}{|T_{\rho}|}\big{(}\#\mathbb{T}^{\prime}[T_{\rho}]\big{)}.$
@@
We conclude that $\mathbb{T}$ satisfies Condition (ii) from Definition 2.15, and hence Conclusion (B) of Proposition 4.1 holds. ∎
5 Sticky Kakeya and the self-similar Convex Wolff Axioms
@@ 5.1 Extremal Kakeya sets satisfy the self-similar Convex Wolff Axioms
Suppose $\omega$ from Definition 3.1 is positive. Then for all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be $\eta$ Assouad-extremal. Then there exists $\mathbb{T}^{\prime}\subset\mathbb{T}$ so that $(\mathbb{T}^{\prime},Y)_{\delta}$ is $\varepsilon$ Assouad-extremal, and $\mathbb{T}^{\prime}$ satisfies the self-similar Convex Wolff axioms with error $\delta^{-\varepsilon}$.
@@
Let $\delta\in(0,\delta_{0}]$ and let $(\mathbb{T},Y)_{\delta}$ be $\eta$ Assouad-extremal. By pigeonholing, we can select a subset $\mathbb{T}_{1}\subset\mathbb{T}$ with #$\#\mathbb{T}_{1}\gtrsim\delta^{\eta}(\#\mathbb{T})$ so that $|Y(T)|\geq\frac{1}{2}\delta^{\eta}|T|$ for each $T\in\mathbb{T}_{1}$. Then #$\#\mathbb{T}_{1}=\delta^{-\beta}$ for some $\beta\geq\alpha-2\eta$ (recall that $\alpha$ is defined in (3.4)), and $\mathbb{T}_{1}$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}$. Furthermore, for all $\rho_{1},r_{1}$ with $\rho_{1}\leq\delta^{\eta}r_{1}$, and all balls $B$ of radius $r_{1}$, we have
@@
(In the above equation, $\omega(\beta+\tau)$ refers to the function $\omega$ evaluated at the (positive) number $\beta+\tau$). Provided $\eta_{1}$ is selected sufficiently small (depending on $\tau$), we have $\beta+\tau\geq\alpha-2\eta_{1}+\tau>\alpha$, and hence by the definition of $\alpha$ from (3.4), we have $\gamma>0$. Hence if $\eta_{1}$ is sufficiently small, then
@@
Step 4. We will show that Conclusion (D) from Proposition 4.1 cannot hold. Suppose to the contrary that Conclusion (D) holds, i.e. there exists $\rho\in[\delta^{1-\varepsilon},1]$, a $\rho$-tube $T_{\rho}$, and a set $\tilde{\mathbb{T}}\subset\mathbb{T}_{1}[T_{\rho}]$ satisfying Items (D.i) and (D.ii) from Conclusion (D). For each $T\in\tilde{\mathbb{T}}$, let $\tilde{Y}(\tilde{T})\subset Y(\tilde{T})$ be the sub-shading produced by Lemma 2.3; this shading is regular at scales $\geq\delta$ and satisfies $|\tilde{Y}(\tilde{T})|\geq\frac{1}{2}|Y(\tilde{T})|$.
Let $\tilde{\delta}=\delta/\rho\in[\delta,\delta^{\varepsilon}]$ and let $\tilde{\mathbb{T}}_{\tilde{\delta}}$ be the unit rescaling of $\tilde{\mathbb{T}}$ relative to $T_{\rho}$, i.e. the tubes in $\tilde{\mathbb{T}}_{\tilde{\delta}}$ are the images of the tubes in $\tilde{\mathbb{T}}$ under the affine map $\phi_{T_{\rho}}$. Let $\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})$ be the image of $\tilde{Y}(\tilde{T})$ under this same rescaling. We have $|\tilde{Y}_{\tilde{\delta}}(\tilde{T})|\gtrsim\delta^{\eta}|\tilde{T}_{\tilde{\delta}}|\geq\tilde{\delta}^{\eta/\varepsilon}|\tilde{T}_{\tilde{\delta}}|$, and we will select $\eta$ and $\delta_{0}$ sufficiently small so that $|\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})|\geq\tilde{\delta}^{\eta_{1}}|\tilde{T}_{\tilde{\delta}}|$ for all $\tilde{T}_{\tilde{\delta}}\in\tilde{\mathbb{T}}_{\tilde{\delta}}$. The tubes in $\tilde{\mathbb{T}}_{\tilde{\delta}}$ are essentially distinct; by Item (D.i) we have #$\#\tilde{\mathbb{T}}_{\tilde{\delta}}\geq\tilde{\delta}^{-\beta-\tau}$; and by Item (D.ii) we have that $\tilde{\mathbb{T}}_{\tilde{\delta}}$ satisfies the Convex Wolff Axioms with error $\lesssim\tilde{\delta}^{-\eta_{1}}$.
@@
$\Big{|}B_{2}\cap N_{\rho_{2}}\Big{(}\bigcup_{\tilde{T}_{\tilde{\delta}}\in\tilde{\mathbb{T}}_{\tilde{\delta}}}\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})\Big{)}\Big{|}\geq(\rho_{2}/r_{2})^{\omega-\gamma/8}|B_{2}|.$ (5.5)
Let $E_{2}=\phi_{T_{\rho}}^{-1}(B_{2})$; this is an ellipsoid of dimensions $\rho r_{2}\times\rho r_{2}\times\rho$. We claim that
$\Big{|}E_{2}\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}\gtrapprox\delta^{-\eta}(\rho_{2}/r_{2})^{\omega-\gamma/8}|E_{2}|.$ (5.6)
To verify this claim, we argue as follows. Let $\tilde{T}_{\tilde{\delta}}\in\tilde{\mathbb{T}}_{\tilde{\delta}}$, let $\tilde{x}\in\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})$, and let $\tilde{B}=B(\tilde{x},\rho_{2})$. Let $E=\phi_{T_{\rho}}^{-1}(\tilde{B})$; this is an ellipsoid of dimensions $\rho_{2}\rho\times\rho_{2}\rho\times\rho_{2}$, and let $x=\phi_{T_{\rho}}^{-1}(\tilde{x})$ be the center of $E$. Similarly, let $\tilde{T}_{\delta}$ and $\tilde{Y}(\tilde{T})$ be the $\delta$ tube and its shading corresponding to $\tilde{T}_{\tilde{\delta}}$ and the shading $\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})$. Then $E\cap\tilde{T}$ contains a tube-segment of length $\geq\rho_{2}/2$ centered at $x$, i.e. $E\cap\tilde{T}\supset B(x,\rho_{2}/2)$. Since $\tilde{Y}(\tilde{T})$ is regular at scales $\geq\delta$ (and $\rho_{2}\geq\tilde{\delta}=\delta/\rho)$, we have that $|E\cap\tilde{Y}(\tilde{T})|\gtrapprox\delta^{\eta}\rho_{2}\delta^{2}$, i.e. $E\cap\tilde{Y}(\tilde{T})$ contains $\gtrapprox\delta^{\eta}(\rho_{2})/(\rho\rho_{2})=\delta^{\eta}\rho^{-1}$ points that are $10\rho\rho_{2}$-separated. For each such point $y$, we have that $B(y,\rho\rho_{2})\subset 2E$, and these balls are disjoint. Since each of these balls has volume roughly $(\rho\rho_{2})^{3}$, we conclude that
$|2E\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}^{\prime}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T}^{\prime})\Big{)}\Big{|}\geq|2E\cap N_{\rho\rho_{2}}(\tilde{Y}(\tilde{T}))|\gtrapprox\delta^{\eta}\rho^{2}\rho_{2}^{3}.$ (5.7)
Next, let $\{\tilde{x}_{i}\}_{i=1}^{N}$ be a $2\rho_{2}$-separated subset of $B_{2}\cap\bigcup_{\tilde{\mathbb{T}}_{\tilde{\delta}}}\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})$ of cardinality
$N\sim\rho_{2}^{-3}\Big{|}B_{2}\cap N_{\rho_{2}}\Big{(}\bigcup_{\tilde{T}_{\tilde{\delta}}\in\tilde{\mathbb{T}}_{\tilde{\delta}}}\tilde{Y}_{\tilde{\delta}}(\tilde{T}_{\tilde{\delta}})\Big{)}\Big{|}\geq\rho_{2}^{-3}(\rho_{2}/r_{2})^{\omega-\gamma/8}|B_{2}|,$
with the property that $B(\tilde{x}_{i},2\rho_{2})\subset B_{2}$ for each index $i$.
For each index $i$, let $E_{i}=\phi_{T_{\rho}}^{-1}(B(\tilde{x}_{i},\rho_{2}));$ this is an ellipsoid of dimensions $\rho_{2}\rho\times\rho_{2}\rho\times\rho_{2}$ (this introduces a slight clash of notation, since we previously introduced the ellipsoid $E_{2}$ of dimensions $\rho r_{2}\times\rho r_{2}\times\rho$). The ellipsoids $\{2E_{i}\}_{i=1}^{N}$ are disjoint and contained in $E_{2}$. Applying (5.7) to each such ellipsoid, we have
$\begin{split}\Big{|}E_{2}\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}&\geq\sum_{i=1}^{N}\Big{|}2E_{i}\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}\\ &\gtrapprox N\delta^{\eta}\rho^{2}\rho_{2}^{3}\\ &\gtrapprox\delta^{\eta}\rho^{2}(\rho_{2}/r_{2})^{\omega-\gamma/8}|B_{2}|\\ &=\delta^{\eta}(\rho_{2}/r_{2})^{\omega-\gamma/8}|E_{2}|.\end{split}$ (5.8)
This establishes (5.6), as claimed.
Next, by (5.6) and pigeonholing we can select a ball $B\subset E_{2}$ of radius $r_{2}\rho$ so that
$\Big{|}B\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}\gtrapprox\delta^{\eta}(\rho_{2}/r_{2})^{\omega-\gamma/8}|B|\geq(\rho_{2}/r_{2})^{\omega-\gamma/8+\eta/(\varepsilon\eta_{1})}|B|.$
If we select $\delta_{0}$ and $\eta$ sufficiently small depending on $\gamma,\varepsilon,$ and $\eta_{1}$, then
$\Big{|}B\cap N_{\rho\rho_{2}}\Big{(}\bigcup_{\tilde{T}\in\tilde{\mathbb{T}}}\tilde{Y}(\tilde{T})\Big{)}\Big{|}\geq 2(\rho_{2}/r_{2})^{\omega-\eta}|B|.$ (5.9)
Finally, observe that the scale separation $\frac{\rho\rho_{2}}{\rho r_{2}}$ satisfies $\frac{\rho\rho_{2}}{\rho r_{2}}=\frac{\rho_{2}}{r_{2}}\leq\rho^{\eta_{1}}\leq\delta^{\eta_{1}\varepsilon}\leq\delta^{\eta}$, provided we select $\eta$ sufficiently small. But then (5.9) contradicts (5.1). Hence Conclusion (D) cannot hold.
@@
It suffices to show that $\omega=0$. Suppose instead that $\omega>0$. Let $\eta_{0},\delta_{0}$ be the output from Theorem 5.2, with $\varepsilon=\omega/3$. By Proposition 5.1, there exists $\delta\in(0,\delta_{0}]$ and a set $(\mathbb{T}^{\prime},Y)_{\delta}$ with discretized Assouad dimension $\leq 3-\omega/3$ that satisfies the self-similar Convex Wolff Axioms (and hence the Convex Wolff Axioms at every scale) with error $\delta^{-\eta_{0}}$, and for which $\{Y(T)\}$ is a $\delta^{\eta_{0}}$ dense shading. But then (5.10) says that $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $33-\omega/$ 3, which is a contradiction. ∎
@@
6 Kakeya estimates for tubes satisfying the Convex Wolff Axioms at every scale